OpenStudy (anonymous):

Another giving another medal before I leave... 4^1+2x=1024

OpenStudy (anonymous):

x=510

OpenStudy (anonymous):

it's 4 to the 1+2x power equals 1024

OpenStudy (anonymous):

akshay how did you get 510?

OpenStudy (anonymous):

\[4^{1+2x}=1024\]

OpenStudy (anonymous):

right if

OpenStudy (anonymous):

@akshay1234 solved other equation

OpenStudy (mathmale):

Please double-check that you've copied this problem correctly. It's highly unusual to see 4^1. Use parentheses to group that 1 with that 2x, or, better yet, use Equation Editor (as you just have). @Opulentgal: Please now try solving this equation for x, showing your work.

OpenStudy (anonymous):

I asked the teacher what it was and she said 4 to the 1+2x power equals 1024

OpenStudy (anonymous):

First of all right 1024 as 2^10 Second of all right :\[4^{1+2x}=2^{2^{1+2x}}=2^{2+4x}\]

OpenStudy (mathmale):

I believe that's the correct interpretation. I'm curious how you'll approach the solution. What would you do first?

OpenStudy (anonymous):

I need to know how to express exponents when I type...

OpenStudy (anonymous):

well mm, I would try to solve for x?

OpenStudy (anonymous):

after this notation your equation looks like: \[2^{10}=2^{2+4x}\]

OpenStudy (mathmale):

That's a separate question, and a good one....but please let's solve \[4^{1+2x}=1024\] first. Based upon the hints given you above, can you now solve for x, @opulentgal?

OpenStudy (anonymous):

express the left side as a log to bring down the exponent?

OpenStudy (mathmale):

@ifrimpanainte: Thanks for your very appropriate guidance.

OpenStudy (anonymous):

If is good! :)

OpenStudy (mathmale):

\[2^{10}=2^{2+4x}\]

OpenStudy (mathmale):

is correct (and yes, "good"). Hint: How would you now simplify \[2^{10}=2^{2+4x}\]???

OpenStudy (anonymous):

cut 4 both side

OpenStudy (mathmale):

@akshay1234: Thanks, but would you please let Opulentgal try to find her own solution?

OpenStudy (anonymous):

1+2x=4 remain

terenzreignz (terenzreignz):

Wrong tag. Anyway, @akshay1234 \[\Large 4^4\] has no part in this :P

OpenStudy (anonymous):

not getting it because since I am methodical, I need step by step; working on this end...

terenzreignz (terenzreignz):

You mentioned logs @Opulentgal Are you more comfortable with those?

terenzreignz (terenzreignz):

@akshay1234 Thank you for that correction. Now, before jumping right into it, first, try to understand what method works for the asker in question here ^_^

OpenStudy (anonymous):

you got 1+2x=5 then send 1 right side

terenzreignz (terenzreignz):

Virtrual <shush> @akshay1234 no spoilers ^_^

OpenStudy (anonymous):

2x=4 comes up then x=4/2

OpenStudy (anonymous):

i need to get the 1024 to match the base on the left side so exponents be = one to one

terenzreignz (terenzreignz):

<sigh>

OpenStudy (anonymous):

x=2

terenzreignz (terenzreignz):

Okay, that's a good way to do it too, @Opulentgal Can you express 1024 as 4 raised to some exponent?

OpenStudy (anonymous):

you can cheak by puting x value in your question ok

OpenStudy (mathmale):

@akshay1234 Would you PLEASE let Opulentgal do her work herself? On OpenStudy we do NOT do other people's work for them nor give them the answers outright. Stop now, please.

terenzreignz (terenzreignz):

I'm sorry, @Opulentgal For two things, firstly, this mix-up concerning akshay And second, there isn't really any methodical way to get what exponent you need to raise 4 to so that you get 1024... Even if you use logs, you still have to know for a fact that 1024 is in fact, 4 raised to...what? ^_^

OpenStudy (mathmale):

Assuming that \[2^{10}=2^{2+4x}\] is correct, please note, Opulentgal, that the 2 sides have the same base. Is there a way to eliminate that common base?

terenzreignz (terenzreignz):

^Or work from there. lol :)

OpenStudy (anonymous):

well mm, you could bring down the exponents since the bases are the same..right

terenzreignz (terenzreignz):

Sometimes, wording is crucial, especially when you're explaining your work. Since the bases are equal, their exponents must also be equal ^is the sound reasoning for this. So, your next step should be...?

OpenStudy (mathmale):

"well mm, you could bring down the exponents since the bases are the same..right?" Yes! "Since the bases are equal, their exponents must also be equal."

OpenStudy (anonymous):

10=2+4x 10-2=2+4x-2 8=4x 8/4=4x/4 2=x

terenzreignz (terenzreignz):

Bravo. ^_^

OpenStudy (anonymous):

thanks mm and terenz, I must give a medal but to whom, wish I could give to both of you...

OpenStudy (mathmale):

Really, really cool, Opulentgal!! Please take notes on this conversation, so that you'll have something helpful to refer to later on.

terenzreignz (terenzreignz):

Don't worry about medals. It's enough that you learned ^_^

OpenStudy (mathmale):

Give terenzreignz the medal and reward me by continuing the positive involvement you've shown here!

terenzreignz (terenzreignz):

Way ahead of you. As usual, a pleasure working with you, professor :) ^_^

OpenStudy (anonymous):

mm,. I pay it forward by answering history questions since I like history! :)

OpenStudy (mathmale):

Great idea!!!